OpenStudy (anonymous):

The sum of the squares of 3 consecutive positive integers is 116. What are the numbers? Which of the following equations is used in the process of solving this problem? 3n2 + 5 = 116 3n2 + 3n + 3 = 116 3n2 + 6n + 5 = 116

OpenStudy (jiteshmeghwal9):

let the smallest of the consecutive integers be 'x'. then medium number=(x+1) largest number=(x+2) A/q, \[(x)^2+(x+1)^2+(x+2)^2=?\]